solve by systems of equations.
THe units digit of a two digit number is twice the tens digit. If the digits are reversed, the new number is 36 more than the origional number. Find the number.
im lost.
show all work please

Answer by stanbon(75887)

THe units digit of a two digit number is twice the tens digit. If the digits are reversed, the new number is 36 more than the origional number. Find the number.
-------------
Let the number be 10t+u
---
Equations:
u = 2t
10u+t = 10t+u+36
-----------------------
Substitute for "u" and solve for "t":
10(2t)+t = 10t+2t+36
21t = 12t+36
9t = 36
t = 4
---
Since u = 2t, u = 8
----
Original number: 10t+u = 48
=================================
cheers,
Stan H.